how do i solve with substitution?
2x+3y=7
4x+3y=13

Answer by reviewermath(1029)

2x + 3y = 7 ---> 3y = 7 - 2x
4x + 3y = 13 ---> 4x + (7 - 2x) = 13
2x + 7 = 13
x = 3
3y = 7 - 2(3)
y = 1/3
Answer: x = 3 and y = 1/3